# Convex Formulation of the Zero Emission Vessel Route Planning Problem

### v2

@@ 7.1 Transportation network
We model sea transportation services in the northern Baltic Sea region. The route network connects five major ports in Sweden, Estonia, the Finnish mainland, and the autonomous region of the Åland Islands. Connections include both short-distance shuttle services and overnight cruises. Most vessels are ro-pax ferries offering transportation for both ro-ro cargo and passengers. The total number of passengers transported in all the routes was 11.8 million in 2024 (AS Tallink Grupp, 2025). Ro-ro cargo volume, consisting of passenger vehicles, trucks, and trailers, was 2.2 million in the same period.
Conventional combustion engine-powered vessels currently operate all the routes in this region. Fast bunkering of liquid hydrocarbon fuels enables short turnaround times. In particular, the 160 nautical mile route from Turku (Finland) to Stockholm (Sweden) (figure 9) runs on a 24-hour periodic schedule with less than an hour turnaround time in both end ports. Battery-electrification of this route is difficult because the current schedule requires charging powers exceeding 100 megawatts or implementing a battery-swapping scheme.
The other routes’ relaxed schedules or short distances make them better suited for electrification. The long-haul overnight routes from Helsinki and Tallinn to Stockholm run on a 48-hour periodic schedule with 7-hour port visits. The flexibility in these routes creates an opportunity to optimize sailing speeds or add a new port call to the service. For example, the Helsinki–Mariehamn–Stockholm vessels can add a 5-hour shuttle service to Tallinn instead of staying berthed continuously for 7 hours at Helsinki.
@@ 7.2 Cases
We formulate and solve six instances of the ZEVRPP in §6, which are labeled as
@@
The letter on the label stands for the fleet type. In the baseline fleet B, the number of vessels and their principal dimensions match the present fleet of the largest ferry operator in the region. Moreover, we assign the services the same frequencies currently used in the routes. The uniform fleet U consists of interchangeable vessels of the same optimized design allocated for all the routes. The mixed fleet M is assembled from vessels optimized for each route.
The number in the label represents the number of distinct routes in the transportation network. Figure 9 illustrates the current network of four routes. The first alternative network in figure 10 routes traffic from Estonia to Sweden via Finland, eliminating one route. The number of routes is further reduced to two and one while preserving a direct connection between all the mainland ports (figures 11 and 12). In all the alternative networks, vessels in service to and from Sweden call at the Åland Islands to benefit from tax exemption, which enables duty-free sales on board.
We use CVXPY (Diamond and Boyd, 2016) modeling language to represent the problem instances and solve them with the primal-dual interior point algorithm implemented in the solver ECOS (Domahidi et al., 2013). The branch and bound extension handles the integer-valued decision variables. The solver obtains the global optimum in less than 10 seconds on a standard desktop computer in all the problem instances. Table 3 shows the demand input data (the same for all problem instances) for every active origin-destination port pair. The other fixed parameter values can be accessed in the implementation code (see §1.4).
Table 3: 48-hour demands between each port-pair $(i,j):i\in\mathcal{N},j\in\mathcal{N},i\neq j$. The data are given in the format passengers (cargo). One unit is a thousand lane-meters or a thousand passengers. The demand is estimated from route-specific traffic volumes and market shares reported in AS Tallink Grupp (2025)
@@
Figure 9: Northern Baltic Sea ferry service route plan with four routes.
Figure 10: Northern Baltic Sea ferry service route plan with three routes.
Figure 11: Northern Baltic Sea ferry service route plan with two routes.
Figure 12: Northern Baltic Sea ferry service route plan with a single route.
@@ 7.3 Results
Table 4: Optimal values of selected fleet design variables. The annual total fleet cost in millions of EUR is displayed in brackets below the case label.
Case Route $N^{\text{rt}}$ $N^{\text{vessel}}$ $L$ $L^{\text{sup}}$ $V^{\text{GT}}$ $E^{\text{batt}}$ $t^{\text{cell}}$
[m] [m] $10^{3}$ [MWh] [a]
B4 12 6 3 212 191 43.6 38.9 2.1
(216.4) 145 1 2 212 191 55.3 82.8 7.6
245 1 2 203 183 50.4 81.6 7.6
345 2 1 212 191 58.9 62.3 2.5
U4 12 6 1 161 130 21.9 101.6 17.0
(114.5) 145 2 1 161 130 21.9 101.6 4.2
245 2 2 161 130 21.9 101.6 4.6
345 2 3 161 130 21.9 101.6 16.5
M4 12 6 1 109 83 11.1 13.1 1.6
(106.6) 145 2 1 200 101 30.3 131.6 4.3
245 2 1 223 202 54.4 163.6 4.5
345 2 3 133 119 18.7 24.8 3.2
M3 12 4 1 96 63 9.9 5.7 3.6
(103.8) 1245 2 2 185 167 32.8 197.1 4.2
345 2 2 141 127 21.8 24.5 3.9
M2 1245 2 2 210 150 36.8 224.2 4.3
(120.7) 345 2 3 143 120 22.0 24.8 3.9
M1 12345 1 14 166 107 20.5 60.3 6.7
Table 5: Optimal values of selected fleet operation variables. Sailing leg speeds are in the order they appear in the outbound voyage. The value in the brackets is the speed of the same sailing leg in the inbound voyage.
Case Route $v_{1}$ $v_{2}$ $v_{3}$ $v_{4}$
[kn] [kn] [kn] [kn]
B4 12 21.6 (21.6) - - -
145 14.9 (15.0) 15.6 (15.3) - -
245 15.1 (15.1) 15.6 (15.5) - -
345 15.7 (15.6) 15.7 (15.8) - -
U4 12 16.0 (16.0) - - -
145 21.2 (21.3) 23.1 (23.7) - -
245 21.2 (21.2) 23.1 (22.8) - -
345 14.3 (14.2) 14.3 (14.2) - -
M4 12 15.1 (15.1) - - -
145 21.8 (21.7) 22.7 (22.8) - -
245 22.3 (22.3) 23.3 (23.2) - -
345 14.2 (14.1) 14.4 (14.4) - -
M3 12 10.0 (10.0) - - -
1245 27.6 (27.2) 25.6 (25.6) 27.2 (27.0) -
345 13.9 (13.9) 14.6 (14.6) - -
M2 1245 27.7 (27.3) 25.9 (25.9) 27.3 (27.1) -
345 13.9 (13.8) 14.5 (14.5) - -
M1 12345 17.5 (17.5) 16.1 (16.1) 17.5 (17.5) 17.5 (17.5)
@@
Tables 4 and 5 shows optimal values of selected fleet design and operation variables, while table 6 shows the optimal port charger powers. Comparing the solutions for cases B4 and M4, we see that an optimized fleet more than halves the fleet cost compared to the present fleet in the same 4-route network. M4 deploys smaller ships to the short routes and a smaller overall fleet size. On the other hand, two routes operate at a higher frequency, enabled by higher speeds and shorter turnaround times. In the Helsinki-Tallinn route, M4 maintains the same frequency as B4, but manages to reduce speed by optimizing the port turnaround times. The vessel deployed to the cargo-heavy Tallinn-Helsinki route features a short superstructure, which saves building cost, electricity, and port charges.
Deploying a uniform fleet to the 4-route network is only 7.4% more expensive than an optimized fleet (cases U4 and M4). This result is surprising considering that the gross tonnage of the largest vessel in M4 is almost five times larger than the gross tonnage of the smallest vessel. In U4, the batteries are vastly oversized in the short routes. However, the oversizing also brings low cycling load and long lifetime.
@@ 8.1 Demand scenarios
Demand for maritime transportation services varies daily, weekly, and monthly. For example, the monthly passenger traffic volume in the northern Baltic Sea is three times greater in the summer holiday season compared to mid-winter (Statistics Finland, 2025). Incorporating demand scenarios into the problem formulation in §6 is straightforward by introducing indexing of the operational decision variables by the demand scenarios. This extension costs many new decision variables and constraints, including the discrete service frequency variables. However, the problem decomposes into smaller scenario-specific subproblems, only coupled by the fleet design variables. Specialized solution methods can leverage this property.
@@ 8.3 Travel time cost
Appendix A Still water loads
@@
Figure 13: Weight, buoyancy forces and the resulting still water loads. The origin of the coordinate system is at the forward perpendicular.
Appendix B Hull girder vertical section modulus
@@
$\tilde{I}$= $\sum_{i}\left(I_{i}+A_{i}(z_{\text{NA}}-\tilde{z}_{\text{NA}})_{i}^{2}\right)=\sum_{i}\left(\frac{w_{i}h_{i}^{3}}{12}+w_{i}h_{i}(z_{\text{NA}}-\tilde{z}_{\text{NA}})_{i}^{2}\right)$ (42)
@@
Table 7: Properties of the midship cross-section structural elements.
Element NN $w$ $h$ AA $z_{\text{NA}}$ $z_{\text{NA}}-\tilde{z}_{\text{NA}}$
Bottom 1 $B$ $3p^{\text{td}}/2$ $3Bp^{\text{td}}/2$ 0 2D/5-2D/5
Inner bottom 1 $B$ $p^{\text{td}}$ $Bp^{\text{td}}$ $D/10$ 3D/10-3D/10
Ro-ro deck 1 $B$ $p^{\text{td}}$ $Bp^{\text{td}}$ $D/2$ $D/10$
Top deck 1 $B$ $p^{\text{td}}$ $Bp^{\text{td}}$ $D$ $3D/5$
Side plate 2 $p^{\text{sp}}$ $D$ $Bp^{\text{td}}/2$ $D/2$ $D/10$
Bulkhead 1 $p^{\text{sp}}$ $D$ $Bp^{\text{td}}$ $D/2$ $D/10$
@@
Appendix C Hull girder shear flow
@@
(Here $G$ is the shear modulus). For the details of this compensation procedure, see §5 in Varsta et al. (2017) or §6.3 in Mittelstedt (2021). The symmetry properties of the hull girder cross-section result in zero shear flows at the half-section vertical centerline of each horizontal element, as illustrated in figure 14.
@@
$q_{E}=-\frac{53BDp^{\text{dp}}}{200\tilde{I}}$
@@
Top deck A B $B/4$ $p^{\text{td}}$ D D 0
Side shell B D $D/2$ $p^{\text{sp}}$ D $D/2$ $q_{A\rightarrow B}$
Ro-ro deck C D $B/4$ $p^{\text{td}}$ $D/2$ $D/2$ 0
Side shell D E D/10 $p^{\text{sp}}$ $D/2$ $2D/5$ $q_{A\rightarrow D}+q_{C\rightarrow D}$
@@
Appendix D Numerical integration of wetted area
@@
$l^{\mathrm{mid}}_{T}=\int_{0}^{B/2}\left\|\begin{bmatrix}-\frac{dz^{\text{mid}}(x)}{dx}\\ 1\end{bmatrix}\right\|\,dx=\int_{0}^{B/2}\sqrt{1+\left(\frac{2}{B}\right)^{2\beta}(T\beta)^{2}x^{2\beta-2}}\,dx,$ (46)

### v3

@@ 7.1 Transportation network
We model sea transportation services in the northern Baltic Sea region. The route network connects five major ports in Sweden, Estonia, the Finnish mainland, and the autonomous region of the Åland Islands. Connections include both short-distance shuttle services and overnight cruises. Most ships are ro-pax ferries offering transportation for both ro-ro cargo and passengers. The total number of passengers transported in all the routes was 11.8 million in 2024 (AS Tallink Grupp, 2025). Ro-ro cargo volume, consisting of passenger vehicles, trucks, and trailers, was 2.2 million in the same period.
Conventional ships currently operate all the routes in this region. Fast bunkering of liquid hydrocarbon fuels enables short turnaround times. In particular, the 160 nautical mile route from Turku (Finland) to Stockholm (Sweden) (figure 10, top left) runs on a 24-hour periodic schedule with less than an hour turnaround time in both end ports. Battery-electrification of this route using the current schedule requires charging powers exceeding 100 megawatts or implementing a battery-swapping scheme.
The other routes’ relaxed schedules or short distances make them better suited for electrification. The long-haul overnight routes from Helsinki and Tallinn to Stockholm run on a 48-hour periodic schedule with 7-hour port visits. The flexibility in these routes creates an opportunity to optimize sailing speeds or add a new port call to the service. For example, the Helsinki–Mariehamn–Stockholm ships can add a 5-hour shuttle service to Tallinn instead of staying berthed continuously for 7 hours at Helsinki.
@@ 7.2 Cases
We formulate and solve six instances of the ESFSMP in §4.5, which are labeled as
@@
The letter on the label stands for the fleet type. In the baseline fleet B, the number of ships and their principal dimensions match the present fleet of the largest ferry operator in the region. Moreover, we assign the services the same frequencies currently used in the routes. The uniform fleet U consists of interchangeable ships of the same optimized design allocated for all the routes. The mixed fleet M is assembled from ships optimized for each route.
The number in the label represents the number of distinct routes in the transportation network. Figure 10 (top left) illustrates the current network of four routes. The first alternative network in figure 10 (top right) routes traffic from Estonia to Sweden via Finland, eliminating one route. The number of routes is further reduced to two and one while preserving a direct connection between all the mainland ports (bottom panels in figure 10). In all the alternative networks, ships in service to and from Sweden call at the Åland Islands to benefit from tax exemption, which enables duty-free sales on board.
Table 7 shows the demand input data (the same for all problem instances) for every active origin-destination port pair. The other fixed parameter values can be accessed in the implementation code. We solve all the problem instances using ECOS with the default settings.
Table 7: 48-hour demands for each active port-pair. The data are given in the format passengers (cargo). One unit is a thousand lane-meters or a thousand passengers. The demand is estimated from route-specific traffic volumes and market shares reported in AS Tallink Grupp (2025)
@@
Figure 10: Four alternative route plans for a northern Baltic Sea ferry service. A route is a sequence of sailing legs denoted by lines of the same style. A complete voyage consists of outbound and inbound sailing along the same set of legs.
@@ 7.3 Results
Table 8: Optimal values of selected fleet design variables. The annual total fleet cost in millions of EUR is displayed in brackets below the case label. Sailing leg speeds are in the order they appear in the outbound voyage. The value in the brackets is the speed of the same sailing leg in the inbound voyage.
Case Route $N^{\text{trip}}$ $N^{\text{ship}}$ $L$ $L^{\text{sup}}$ $V^{\text{GT}}$ $E^{\text{batt}}$ $t^{\text{cell}}$ $v_{1}$ $v_{2}$ $v_{3}$ $v_{4}$
[m] [m] $10^{3}$ [MWh] [a] [kn] [kn] [kn] [kn]
B4 12 18 3 212 191 43.6 38.9 2.1 21.6 (21.6) - - -
(216.4) 145 2 2 212 191 55.3 82.8 7.6 14.9 (15.0) 15.6 (15.3) - -
245 2 2 203 183 50.4 81.6 7.6 15.1 (15.1) 15.6 (15.5) - -
345 2 1 212 191 58.9 62.3 2.5 15.7 (15.6) 15.7 (15.8) - -
U4 12 6 1 161 130 21.9 101.6 17.0 16.0 (16.0) - - -
(114.5) 145 2 1 161 130 21.9 101.6 4.2 21.2 (21.3) 23.1 (23.7) - -
245 4 2 161 130 21.9 101.6 4.6 21.2 (21.2) 23.1 (22.8) - -
345 6 3 161 130 21.9 101.6 16.5 14.3 (14.2) 14.3 (14.2) - -
M4 12 6 1 109 83 11.1 13.1 1.6 15.1 (15.1) - - -
(106.6) 145 2 1 200 101 30.3 131.6 4.3 21.8 (21.7) 22.7 (22.8) - -
245 2 1 223 202 54.4 163.6 4.5 22.3 (22.3) 23.3 (23.2) - -
345 6 3 133 119 18.7 24.8 3.2 14.2 (14.1) 14.4 (14.4) - -
M3 12 4 1 96 63 9.9 5.7 3.6 10.0 (10.0) - - -
(103.8) 1245 4 2 185 167 32.8 197.1 4.2 27.6 (27.2) 25.6 (25.6) 27.2 (27.0) -
345 4 2 141 127 21.8 24.5 3.9 13.9 (13.9) 14.6 (14.6) - -
M2 1245 4 2 210 150 36.8 224.2 4.3 27.7 (27.3) 25.9 (25.9) 27.3 (27.1) -
(120.7) 345 6 3 143 120 22.0 24.8 3.9 13.9 (13.8) 14.5 (14.5) - -
M1 12345 14 14 166 107 20.5 60.3 6.7 17.5 (17.5) 16.1 (16.1) 17.5 (17.5) 17.5
(193.8) (17.5)
@@
Table 8 shows optimal values of selected fleet design and operation variables, while table 9 shows the optimal port charger powers. Comparing the solutions for cases B4 and M4, an optimized fleet more than halves the fleet cost compared to the present fleet in the same 4-route network. M4 deploys smaller ships to the short routes and a smaller overall fleet size. On the other hand, two routes operate at a higher frequency, enabled by higher speeds and shorter turnaround times. In the Helsinki-Tallinn route, M4 maintains the same frequency as B4, but manages to reduce speed by optimizing the port turnaround times. The ship deployed to the cargo-heavy Tallinn-Helsinki route features a short superstructure, which saves building cost, electricity, and port charges.
Deploying a uniform fleet to the 4-route network is only 7.4% more expensive than an optimized fleet (cases U4 and M4). This result is surprising considering that the gross tonnage of the largest ship in M4 is almost five times larger than the gross tonnage of the smallest ship. In U4, the batteries are vastly oversized in the short routes. However, the oversizing also brings low cycling load and long lifetime.
@@
7.4 Parameter sensitivity
Figure 11 reports the tractional sensitivity of the objective function to fractional changes to selected fixed parameters for the case M3. A sensitivity of +1 means that perturbing the parameter value by 1% increases the objective value by 1% in the vicinity of the optimal solution. The sensitivities are practically useful for guiding the planning process to focus on the most important strategic decisions.
Figure 11: Sensitivity of selected parameters for the case M3.
All the parameter sensitivities are nonzero. Therefore, at least one constraint that includes a given parameter is active at the optimum, governing the solution. Interestingly, the two most sensitive parameters (crew cost and port charge) are unrelated to the battery-electric powertrain. The sensitivity of battery density is positive, because battery mass per unit of energy increases linearly with density.
@@ 8.1 Demand scenarios
Demand for maritime transportation services varies daily, weekly, and monthly. For example, the monthly passenger traffic volume in the northern Baltic Sea is three times greater in the summer holiday season compared to mid-winter (Statistics Finland, 2025). Incorporating demand scenarios into the problem formulation in §4.5 is straightforward by introducing indexing of the operational decision variables by the demand scenarios. This extension costs many new decision variables and constraints, including the discrete service frequency variables. However, the problem decomposes into smaller scenario-specific subproblems, only coupled by the fleet design variables. Specialized solution methods can leverage this property.
@@ 8.3 Travel time cost
9 Conclusions
The design of a battery-electric liner shipping service presents a complex optimization problem due to the strong coupling of decisions concerning cargo routing, sailing schedules, charger powers, and the ship sizing subdomains. Many ship owners will need to deploy battery-electric fleets in the near future, as increasingly strict emission regulations are enforced. The ro-ro, ro-pax, and container feeder segments constitute over 1,350 ships in total in European Union waters alone (European Parliament, 2025). As the transition to an electric marine transportation system involves a clean-sheet design of all aspects of the system, there is a unique opportunity to apply rigorous mathematical optimization techniques to make a large impact.
This paper has presented a convex formulation of the battery-electric liner shipping fleet size and mix problem. The problem is to minimize total fleet purchase and operation cost while modeling everything from fleet size down to the hull girder plate thickness of each ship. A typical real-world numerical case is represented by a network consisting of six ports, four services, and 30 origin-destination demand arcs. The solve time is less than two minutes on a desktop computer, demonstrating the novel possibility for fast fleet design optimization by employing computationally efficient convex optimization methods.
Acknowledgement
This work was supported by the Future Shipping Electrified (FUSE) research project through Business Finland under grant 3545/31/2023.
Appendix A Changes of variables
A nonlinear optimization problem defined by log-convex posynomial functions is written as
minimize $\sum_{k=1}^{K_{0}}c_{k}\prod_{l=1}^{n}x_{l}^{a_{lk}}$ (52)
subject to $\sum_{k=1}^{K_{m}}c_{k}\prod_{l=1}^{n}x_{l}^{a_{lk}}\leq 1,\quad m=1,.,M.$
In the standard form (52), the objective is a nonconvex function and the constraints define nonconvex sets. However, the problem has a special structure that can be exploited. A logarithmic change of variables $y_{j}=\log x_{j}$ (such that $x_{l}=\exp y_{l}$) allows a transformation that convexifies (52). Specifically, the log of a posynomial is convex in $y$:
$\log\sum_{k=1}^{K}c_{k}\prod_{l=1}^{n}x_{l}^{a_{lk}}=\log\sum_{k=1}^{K}c_{k}\exp{(a_{k}^{T}y)}$ (53)
$=\log\sum_{k=1}^{K}\exp{(\log c_{k}+a_{k}^{T}y)}$
because log-sum-exp is a convex function (Boyd and Vandenberghe, 2004) and the argument of the exponential function is linear in $y$.
Using the log transformed posynomials of the form (53), and the abbreviated constant $b_{k}=\log c_{k}$, the log-convex problem is reformulated as the convex optimization problem
minimize $\log\sum_{k=1}^{K_{0}}\exp{(b_{0k}+a_{0k}^{T}y)}$ (54)
subject to $\log\sum_{k=1}^{K_{m}}\exp{(b_{mk}+a_{mk}^{T}y)}\leq 0,\quad m=1,.,M.$
The optimal solution of the original nonconvex problem (52) is easily recovered from a solution of the transformed problem (53) since $x_{l}^{*}=\exp y_{l}^{*}$.
Figure 12: Feasible set of a log-convex optimization problem with two variables. The top panel shows the original nonconvex problem, and the bottom panel shows the log-transformed convex problem. The solid red cross is the optimal solution.
minimize $x_{1}^{-1}$
[…]
$x_{2}\leq 2x_{1}^{-1},\,x_{2}\leq 3.$
Appendix B Still water loads
@@
Figure 13: Weight, buoyancy forces, and the resulting still water loads. The origin of the coordinate system is at the forward perpendicular.
Appendix C Hull girder section modulus
@@
$\tilde{I}$= $\sum_{i}\left(I_{i}+A_{i}(z_{\text{NA}}-\tilde{z}_{\text{NA}})_{i}^{2}\right)$ (55)
$=$ $\sum_{i}\left(\frac{w_{i}h_{i}^{3}}{12}+w_{i}h_{i}(z_{\text{NA}}-\tilde{z}_{\text{NA}})_{i}^{2}\right)$
@@
Table 10: Properties of the midship structural elements.
Element NN $w$ $h$ AA $z_{\text{NA}}$ $z_{\text{NA}}$
$\tilde{z}_{\text{NA}}$
Bottom 1 $B$ $\frac{3p^{\text{td}}}{2}$ $\frac{3Bp^{\text{td}}}{2}$ 0 $\frac{-2D}{5}$
Inner bottom 1 $B$ $p^{\text{td}}$ $Bp^{\text{td}}$ $\frac{D}{10}$ $\frac{-3D}{10}$
Ro-ro deck 1 $B$ $p^{\text{td}}$ $Bp^{\text{td}}$ $\frac{D}{2}$ $\frac{D}{10}$
Top deck 1 $B$ $p^{\text{td}}$ $Bp^{\text{td}}$ $D$ $\frac{3D}{5}$
Side plate 2 $p^{\text{sp}}$ $D$ $\frac{Bp^{\text{td}}}{2}$ $\frac{D}{2}$ $\frac{D}{10}$
Bulkhead 1 $p^{\text{sp}}$ $D$ $Bp^{\text{td}}$ $\frac{D}{2}$ $\frac{D}{10}$
@@
Appendix D Hull girder shear flow
@@
(Here $G$ is the shear modulus.) For the details of this compensation procedure, see §5 in Varsta et al. (2017) or §6.3 in Mittelstedt (2021). The symmetry properties of the hull girder cross-section result in zero shear flows at the half-section vertical centerline of each horizontal element, as illustrated in figure 14.
@@
$q_{E}=-\frac{53BDp^{\text{td}}}{200\tilde{I}}$
@@
Top deck A B $\frac{B}{4}$ $p^{\text{td}}$ D D 0
Side shell B D $\frac{D}{2}$ $p^{\text{sp}}$ D $\frac{D}{2}$ $q_{A\rightarrow B}$
Ro-ro deck C D $\frac{B}{4}$ $p^{\text{td}}$ $\frac{D}{2}$ $\frac{D}{2}$ 0
Side shell D E $\frac{D}{10}$ $p^{\text{sp}}$ $\frac{D}{2}$ $\frac{2D}{5}$ $q_{A\rightarrow D}$
$q_{C\rightarrow D}$
@@
Appendix E Estimation of wetted area
@@
$l^{\mathrm{mid}}_{T}$ $=\int_{0}^{B/2}\left\|\begin{bmatrix}-\frac{dz^{\text{mid}}(x)}{dx}\\ 1\end{bmatrix}\right\|\,dx$ (59)
$=\int_{0}^{B/2}\sqrt{1+\left(\frac{2}{B}\right)^{2\beta}(T\beta)^{2}x^{2\beta-2}}\,dx,$
@@
Appendix F Linear problem formulation
We need two new sets for the mathematical description of the linear problem: let $\mathcal{U}$ be the set of ship types indexed by $u$ and $\mathcal{V}_{u}$ be the set of discrete speed levels for ship type $u$ indexed by $v$.
The energy consumption of a ship type $u$ sailing at speed $v$ from port $i$ to $j$ is a fixed parameter $E_{ijvu}$. Ship type $u$ has capacity $q^{\text{cap}}_{cu}$ of cargo $c$. The ship purchase cost $C^{\text{ship}}_{u}$ and port charge $C^{\text{port}}_{iu}$ are also ship type dependent. The transit time from port $i$ to $j$ depends on the speed level $v$ and is represented by $t^{\text{transit}}_{ijv}$. The cargo handling time $t^{\text{cargo}}_{ij}$ is fixed at the destination port $j$. We use binary variables to select one frequency per service from available frequencies represented by $N^{\text{trip}}_{f}$.
We define the following new decision variables: $x_{ijsvu}\in[0,1]$ is a weight for speed level $v$ for ship type $u$ on service $s$ on leg $(i,j)$. As the weights sum to one on each leg, service, and ship type, the true speed is a linear combination of the discrete speed levels. The charging time at the destination port of leg $(i,j)$ in service $s$ for ship type $u$ is $t^{\text{cha}}_{ijsu}\in\mathbb{R}_{++}$ and the total roundtrip duration in service $s$ is $t^{\text{trip}}_{s}\in\mathbb{R}_{++}$. The ship types and service frequencies are selected using binary variables. Let $z_{fs}\in\{0,1\}$ equal 1 if frequency $f$ is chosen on service $s$ and 0 otherwise and similarly let $y_{us}\in\{0,1\}$ equal 1 if ship type $u$ is selected and 0 otherwise. The nonnegative discrete variable $N^{\text{ship}}_{su}\in\mathbb{Z}_{\geq 1}$ is the number of ships of type $u$ deployed on service $s$. In addition, we need auxiliary variables $q_{ijcs}^{\text{logapp}}\in\mathbb{R}$ for the piecewise linear approximation of the log utility function of $q_{ijcs}$.
The problem formulation is as follows: Some text here that flows to the next line a bit more minimize
$\sum_{s\in\mathcal{S}}\sum_{u\in\mathcal{U}}\Biggl[C^{\text{ship}}_{u}N^{\text{ship}}_{su}+\sum_{(i,j)\in\mathcal{L}_{s}}\sum_{f\in\mathcal{F}}N^{\text{trip}}_{f}z_{sfu}C_{ju}^{\text{port}}$ (63)
$+\sum_{(i,j)\in\mathcal{L}_{s}}\sum_{f\in\mathcal{F}}C_{j}^{\text{el}}N^{\text{trip}}_{f}z_{sfu}t^{\text{cha}}_{ijs}P_{j}\Biggr],$
subject to
$N^{\text{ship}}_{su}\leq N^{\text{ship,max}}_{u}y_{us},\quad s\in\mathcal{S},\,u\in\mathcal{U},$ (64)
$\sum_{u\in\mathcal{U}}y_{us}=1,\quad s\in\mathcal{S},$ (65)
$\sum_{f\in\mathcal{F}}z_{sfu}=y_{us},\quad s\in\mathcal{S},\,u\in\mathcal{U},$ (66)
$\begin{array}[]{lr}t^{\text{trip}}_{s}=\sum_{(i,j)\in\mathcal{L}_{s}}\Bigl(\sum_{v\in\mathcal{V}_{u}}\sum_{u\in\mathcal{U}}t^{\text{transit}}_{ijv}x_{ijsvu}\\ +t^{\text{port}}_{ijs}\Bigr),\quad s\in\mathcal{S},\end{array}$ (69)
$\sum_{u\in\mathcal{U}}\sum_{f\in\mathcal{F}}N^{\text{trip}}_{f}z_{sfu}t_{s}^{\text{trip}}\leq\sum_{u\in\mathcal{U}}N^{\text{ship}}_{su}t_{s}^{\text{route}},\quad s\in\mathcal{S},$ (70)
$t^{\text{cha}}_{ijs}=\frac{1}{P_{j}}\sum_{v\in\mathcal{V}_{u}}\sum_{u\in\mathcal{U}}E_{ijvu}x_{ijsvu},\,(i,j)\in\mathcal{L}_{s},\,s\in\mathcal{S},$ (71)
$t_{ijs}^{\text{port}}\geq\max\{t_{ij}^{\text{cargo}},t_{ijs}^{\text{cha}}\},\quad(i,j)\in\mathcal{L}_{s},\,s\in\mathcal{S},$ (72)
$\sum_{v\in\mathcal{V}_{u}}x_{ijsvu}=y_{us},\quad(i,j)\in\mathcal{L}_{s},\,s\in\mathcal{S},\,u\in\mathcal{U},$ (73)
$\sum_{v\in\mathcal{V}_{u}}E_{ijvu}x_{ijsvu}\leq E^{\text{max}}_{u}y_{us},\begin{array}[]{ll}&(i,j)\in\mathcal{L}_{s},\\ &s\in\mathcal{S},\,u\in\mathcal{U},\end{array}$ (76)
[…]
$\sum_{s:(i,j)\in\mathcal{A}_{s}}q_{ijcs}\leq q_{ijc}^{\text{dem}},\quad c\in\mathcal{C},\ (i,j)\in\mathcal{A},$ (83)
$\sum_{(i,j)\in\mathcal{A}}\sum_{s:(i,j)\in\mathcal{A}_{s}}\sum_{c\in\mathcal{C}}\alpha_{ijc}q^{\text{logapp}}_{ijcs}\geq U^{\text{min}},$ (84)
$\begin{array}[]{lr}q^{\text{logapp}}_{ijcs}\leq 2\log(\hat{n})q_{ijcs}+\log(n)+2n\log(1/\hat{n}),\\ \quad(i,j)\in\mathcal{L}_{s},\,c\in\mathcal{C},\,s\in\mathcal{S},\,n=1/2,1,2,4,\ldots\end{array}$ (87)
where $\hat{n}=(n+1/2)/n$.
The objective function (63) minimizes the total annualized cost. The three terms are the ship purchase and running cost, port charges, and electricity cost. In contrast to §4.5, the shore charger installation cost is excluded because charger powers are fixed.
The constraints are as follows:
  1. 1.
    Constraints (64)-(66) encode logic relations: a nonzero number of ships of a given type can be deployed on a service only if the ship type is selected. Each service uses only one ship type and one frequency.
  2. 2.
    Constraints (69)-(73) couple service frequency and time spent on transit, port cargo handling, and charging. The total round-trip time (69) is the sum of speed-dependent sailing times and port turnaround times over all the legs. Ships must complete all their round-trips within the available time of the planning horizon as enforced by (70). The charging time at the destination port of a leg is a function of the energy consumption on the leg by equation (71). The port operation that takes the longest determines the turnaround time by (72). The speed weights of the selected ship type sum to one in the equation (73).
  3. 3.
    The inequality (76) states that the energy consumption on any leg cannot exceed the battery capacity.
  4. 4.
    The capacity constraints (79)-(82) ensure that the sum of cargo quantity that flows along a sailing leg does not exceed ship capacity. The constraints for the outbound and inbound parts of the voyage are given separately. The demand constraints (83) state that the total cargo quantity transported from an origin port to a destination port by all the services does not exceed the demand. The inequality formulation implies that demand can be rejected.
  5. 5.
    The transportation service must provide a minimum level of utility by (84). The constraints (87) encode a piecewise linear approximation of the nonlinear log utility function.
Lastly, we reformulate the expression $z_{sfu}t^{\text{cha}}_{ijs}$ in the objective (63) and the expression $z_{sfu}t_{s}^{\text{trip}}$ in the constraints (69). These terms consist of products of a binary variable and a nonnegative continuous variable. The idea is to replace the bilinear product term with a new continuous variable that is forced to take a positive value only when the binary variable is one.
To reformulate the product term in (63), we introduce new auxiliary variables $\tilde{t}_{sfu}^{\text{cha}}\in\mathbb{R}_{+}$ (for all $s\in\mathcal{S},\,f\in\mathcal{F},\,u\in\mathcal{U}$), and the following new constraints:
$\tilde{t}_{sfu}^{\text{cha}}\leq\sum_{(i,j)\in\mathcal{L}_{s}}t_{ijs}^{\text{cha}}-M^{(1)}(1-z_{sfu}),\quad s\in\mathcal{S},\,f\in\mathcal{F},\,u\in\mathcal{U}.$ (88)
Here $M^{(1)}$ is a sufficiently large constant. The last term in (63) is then rewritten linearly as
$\sum_{f\in\mathcal{F}}C_{j}^{\text{el}}N^{\text{trip}}_{f}\tilde{t}_{sfu}^{\text{cha}}P_{j}.$
The constraint (88) states that the left-hand side must be equal to or greater than the sum on the right-hand side if the binary variable is one. Otherwise, the left-hand side is reduced to the lower bound of zero.
The product terms in (69) are reformulated using the same technique as above.